A monopolist faces two totally separated markets with inverse demand p=100 – qA and p=160−2qB respectively. The monopolist has no fixed costs and a marginal cost given by mc= 2 /3 q Find the profit maximizing total output and how much of it that is sold on market A and market B respectively if the monopoly uses third degree price discrimination.
1
Expert's answer
2020-06-09T16:23:11-0400

Marginal Costs (MC) = "\frac {2}{3}q"

Market 1 Inverse demand fxn: P=100 – qA

Total Revenue (TR) = P*Q

TR = "(100 – Q) \times Q"

TR = "100Q \u2013 Q^2"

MR = "\varDelta TR / \varDelta Q = (100Q – Q^2)/ \varDelta Q"

MR = 100 – 2Q

Market 2 Inverse demand fxn: p=160−2qB

Total Revenue (TR) = "P\times Q"

TR = "(160 – 2Q) \times Q"

TR = "160Q \u2013 2Q^2"

MR = "\varDelta TR / \varDelta Q = (160Q – 2Q^2)/ \varDelta Q"

MR = 160 – 4Q

Equilibrium

MR = MC

Market 1: "100 – 2Q = \frac{2}{3}Q"

"\frac{8}{3}Q = 100"

Q = "100\times \frac{3}{8}"

"Q_{1} = 37.5"

P = 100 – Q

P = 100 – 21.43

P = 78.57

TR = "100Q \u2013 Q^2"

TR = "100(37.5) \u2013 37.5^2"

TR = "160(37.5) \u2013 2(37.5^2)"

TR = 2,343.75

Market 2: 160 – 4Q = 2/3Q

14/3Q = 100

Q = "100\times \frac{3}{14}"

"Q_{2} = 21.43"

P = 160 – 2Q

P = "160 – 2\times21.43"

P = 160 – 42.86

P = 117.14

TR = "160Q \u2013 2Q^2"

TR = "160(21.43) \u2013 2(21.43^2)"

TR = 3,428.8 – 918.4898

TR = 2,510.31